OpenStudy (anonymous):

change the cartesian integral into an equivalent polar integral:

OpenStudy (anonymous):

\[\int\limits_{\sqrt{2}}^{2}\int\limits_{\sqrt{4-y^{2}}}^{y} dx dy\]

OpenStudy (amistre64):

i think this is what it looks like at least

OpenStudy (amistre64):

remember that x=rcos(t) y=rsin(t) and r=sqrt(x^2+y^2), but thats really already given as 2

OpenStudy (anonymous):

I think I understand how to get pi/4 < theta < pi/2 But finding the bounds for r is throwing me off.

OpenStudy (amistre64):

well, since r only really changes from 0 to 2; the bounds for r are 0 to 2

OpenStudy (amistre64):

it r changed according to some function of t the r would be bound from 0 to f(t)

OpenStudy (anonymous):

according to the answers, r is bounded by 2 < r < 2csc(theta)

OpenStudy (amistre64):



OpenStudy (anonymous):

but at pi/2 r can be 2 and rad(2)

OpenStudy (amistre64):

it might help to see the limits in this fashion to be able to see how to change them out \[\Large \int_{y=\sqrt{2}}^{y=2}~~~\int_{x=\sqrt{4-y^{2}}}^{x=y} ~~~dx~dy\]

OpenStudy (anonymous):

my book says to plug in x=rcos(theta) etc and magically gets the polar bounds. They dont show any intermediate steps so im really confused.

OpenStudy (amistre64):

\[x= \sqrt{r^2-y^2}\]\[2cos(t) = \sqrt{2^2-(2sin(t))^2}\]\[4cos^2(t) = 4-4sin^2(t)\]\[cos^2(t) +sin^2(t)=1~\text{for any t}\] \[x= y\]\[2cos(t) = 2sin(t)\]\[t=\frac{pi}4\] something along those lines i believe

OpenStudy (amistre64):

y = 2 2sin(t) = 2 sin(t) = 1; when t=pi/2 y=sqrt(2) 2sin(t)=sqrt(2) sin(t)=sqrt(2)/2; when t=pi/4